OpenStudy (anonymous):

what is the measure of angle ACB angle A=20 degrees angle B=70

jhonyy9 (jhonyy9):

so angle A + angle B = 90 the sum of the angles of a triangle is equal 180 degrees this mean that angle ACB what mean angle C will be equal 180 -(70+20) = =180 - 90 = 90 hope this will help you understanding